# A Demonstration of Slowed Electron Drift for PTOLEMY

### v2

@@ Abstract
The novel transverse electromagnetic filtering system proposed by the PTOLEMY collaboration necessitates the precise control of -decay electron motion via drift. In particular, drift speed must be slowed substantially to contain electrons within the radio-frequency antenna region and accurately tune voltage potentials within the filter. This paper presents the experimental design to vary drift speed of -decay electrons using a custom electrode field cage situated between the pole faces of an electromagnet. Matching our results with high-fidelity simulation, we deduce a capacity to increase particle time of flight by a factor of 5 in the field cage’s slow drift region. Limited only by the dimensions of our system, we assert drift speed can be arbitrarily slowed to meet the needs of PTOLEMY’s future detector.
@@ 1 Introduction
1.1 Neutrino Mass Experiments and Status
Juxtaposed to the Standard Model of particle physics, the discovery of neutrino oscillations [1, 2] necessitates a nonzero effective neutrino mass , though it has yet to be precisely measured.111, where and denote neutrino mass eigenstates and electronic mixing matrix elements, respectively. The current model-independent upper limit on is 0.45 eV (90% confidence level), as set by the Karlsruhe Tritium Neutrino (KATRIN) experiment [3]. However, KATRIN and similar molecular tritium experiments are limited to a sensitivity of eV due to broadening caused by internal molecular motion [4, 5] and background from surface radioimpurities [6]. Instead using atomic tritium , recently proposed projects such as PTOLEMY suggest that advances in our understanding of neutrino physics and techniques for high-sensitivity instrumentation could significantly improve sensitivity to measuring [7, 8].
@@ 1.2 The PTOLEMY Experiment
PTOLEMY aims to create a scalable neutrino mass measurement experiment that performs precision spectroscopy of electrons produced near the endpoint energy from the -decay of (equation 1.1). With strong background suppression and high energy resolution, distortions of the emission spectrum endpoint would enable a direct measurement of . Tritium was selected for multiple reasons, including long lifetime and low -value ( keV) [9]. The proposed target substrate is graphene [10], which can efficiently store atomic tritium by binding it to carbon atoms. Extensive work is currently being undertaken by the collaboration and community at large to study the solid state effects on the emission spectra [11, 12, 13, 14].
To suppress background, PTOLEMY employs a technique called cyclotron radiation emission spectroscopy (CRES), which was pioneered by the Project 8 collaboration [15, 16]. Using radio-frequency (RF) antennas, an electron’s kinetic energy is inferred by the relativistic frequency shift of the emitted cyclotron radiation. In the case of a near-endpoint electron, this information is used to guide the particle toward the calorimeter region. Provided enough time to obtain a large cyclotron radiation signal, CRES has been shown to be highly precise and stringently limit backgrounds near the tritium endpoint [17], enabling the possibility of direct differential energy measurement in time coincidence following a non-destructive CRES RF tag.
@@
To attain high measurement resolution, PTOLEMY utilizes transition-edge sensor (TES) microcalorimeters, which precisely measure energy via the strongly temperature-dependent resistance of a cryogenic superconducting material’s phase transition. In order for tritium endpoint electrons to fall within the dynamical range of the TES, static electromagnetic fields are used to reduce their kinetic energy from 18.6 keV to roughly 10 eV. This is performed by PTOLEMY’s novel transverse drift filter [18]. The filter uniquely orients both the magnetic field gradient and electric field orthogonal to the magnetic field, resulting in components of the electron’s motion driven by and . This allows for a compact, scalable detector, with each dimension on the order of 1 meter [19]. In total, this technique could achieve an energy resolution as small as 50 meV [20], opening the possibility to a detection of the Cosmic Neutrino Background [21, 9] in addition to an effective neutrino mass measurement. A schematic of the PTOLEMY prototype is shown in Figure 1.
@@ 1.3 Transport
drift has been extensively researched in the contexts of plasma dynamics, propulsion, and astrophysical phenomena (for instance: [22], [23], and [24]), though is seldom employed deliberately for charged particle transport.222For a particle exposed to an electric field while undergoing cyclotron motion due to a magnetic field (satisfying ), the drift velocity of its guiding center (the center point of circular cyclotron motion) is given as and commonly referred to as drift. While there are studies involving the use of drift to control low-energy positrons [25], contain secondary electrons [26], and create an angular-selective calibration source [27], PTOLEMY’s concept of slow drift for low-energy electron transport is entirely novel. Tethered to the strength of , an electron’s drift velocity can easily be changed by adjusting voltage potentials throughout the system. This feature is crucial to precisely reduce drift velocity near the RF antennas such that there is enough time (on the order of 100 s [28]) to obtain a large enough signal to accurately tag electrons. This paper presents the design and characterization of a setup to demonstrate such slow drift of -decay electrons.
@@ 2.1 Concept
The conceptual setup for the demonstrator is shown in Figure 2. Eight electrode plates are situated between the pole faces of a magnet. Two of these electrodes, the bouncing electrodes, are oriented parallel to the magnet pole faces and supplied with a negative voltage to create a potential barrier to contain electrons within the apparatus (such that the electron motion parallel to the magnetic field exhibits a bouncing behavior), similar to the electrostatic mirror described in [29]. The remaining six are organized into three pairs of vertically offset drift electrodes, with voltages applied to generate an electric field orthogonal to the magnetic field. The potential difference between the center pair of plates is set to be smaller than that of the outer pairs, thereby creating a region of slower drift. In total, an electron injected through one end of the demonstrator undergoes cyclotron motion, bounces between bouncing electrodes, and drifts toward the other end. An example of such a trajectory is shown from a top-down orientation in Figure 3. The increased density of trajectory lines in the center region indicates slower drift. This Figure and all subsequent simulation estimates are generated using CST Studio Suite by Dassault Systèmes [30].
@@ 2.4 PIN Diode Detector
For electron identification, the system size and operating environment necessitate the use of a detector that is both compact and resilient to high magnetic fields. Inspired by the utilization of PIN diodes for low-energy electron detection by KATRIN [32], the OSRAM BPX 61 Silicon PIN diode was selected as a detector for this study. The specific model was chosen for its commercial availability and comparatively low cost. PIN diodes are electrical components containing an undoped intrinsic semiconductor region between a p-type and n-type semiconductor. When operated in reverse bias, they are sensitive to energy depositions of charged particles and light. The BPX 61 additionally features a low dark current of 2 nA and a relatively large active detection area (), limiting electrical noise while increasing possible event count. However, commercial diodes — including the BPX 61 — are typically hermetically sealed by plastic or glass to prevent damage to the semiconductor region. Such covers, though providing little impedance to light, have a great enough stopping power to entirely prevent low-energy electron detection. For our study, the cover was carefully removed and the exterior casing was flattened to the base of the diode, improving sensitivity to higher-energy electrons with larger cyclotron radii. A heat shrink tube was attached around the neck of the diode to prevent torsion/bending from magnetization. An image of the BPX 61 diode and an example of the performed alterations are shown in Figures 7.
Energy threshold characterization of the device was performed using a beta spectrometer (details provided in Supplementary Material). In summary, the spectrometer sequentially selects energy range windows of transmittable electrons, and the PIN diode records counts of events during each iteration. Events are considered to be waveforms whose peak height exceeds a value of 45 mV above baseline after amplification (the electronic and amplification chain are identical to those discussed in Section 2.5). Based on the phase space of allowed particle trajectories in the system (discussed in Section 5), we situate the diode at a shallow angle () with respect to the instrument aperture. In Figure 8, we present a relative trigger efficiency curve for the device obtained from beta spectrometer counts, observing a half-maximal trigger efficiency occurring at keV. To clarify, this result does not reflect the true sensitivity of the device, but rather the trigger efficiency at a threshold safely above the noise floor of our subsequent data-taking conditions. Following manufacturer specifications, we assign an energy resolution of 11%.

### v6

@@ Abstract
To resolve the effective neutrino mass with an energy resolution of 50 meV, the PTOLEMY experiment has proposed a novel transverse electromagnetic filtering process. Substantially reducing the kinetic energy of tritium -decay electrons by counteracting motion from and drift, the PTOLEMY filter requires an input of emitted electron kinematic information to generate a tailored, suitable electric field for each candidate. The collaboration proposes to extract these quantities by using antennae to observe the relativistic frequency shift of emitted cyclotron radiation as an electron transits by drift through a uniform magnetic field region preceding the filter. Electrons must be contained within this region long enough such that an adequate integrated radiated power signal is received to accurately estimate these kinematics. This necessitates a controlled, slowed drift speed. This paper presents the experimental design to vary drift speed of -decay electrons using a custom electrode field cage situated between the pole faces of an electromagnet. Matching our results with high-fidelity simulation, we deduce a capacity to increase particle time of flight by a factor of 5 in the field cage’s slow drift region. Limited only by the dimensions of our system, we assert drift speed can be arbitrarily slowed to meet the needs of PTOLEMY’s future detector. Actualizing such a system is a crucial milestone in developing the detector, enabling future cyclotron radiation measurements, filter implementation, and source injection.
@@ 1 Introduction
1.1 Neutrino Mass Landscape
Juxtaposed to the Standard Model of particle physics, the discovery of neutrino oscillations (between flavor states and ) [1, 2] necessitates that at least two neutrinos are massive particles. Each neutrino flavor thus consists of a superposition of three mass eigenstates, where the mixing components are elements of the 3 3 unitary Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix [3, 4].111For PMNS matrix element , denoting flavor eigenstate, and denoting mass eigenstate. While the magnitudes of the squared differences between the three generations of neutrino mass eigenstates are known with high precision [5], the absolute mass scale underpinning these values has yet to be measured.222It additionally is unknown exactly what the ordering of these mass eigenstates (i.e., the neutrino mass hierarchy) is. The two possible cases, where the third mass eigenstate is the heaviest or lightest of the three, are respectively referred to as Normal Ordering (NO) and Inverted Ordering (IO). Neutrinos, in fact, are the only fundamental fermions whose masses remain unknown. Constraints nonetheless have been established on mass-relevant parameters from cosmological observations and various experiments.
As the most abundant massive particle, neutrinos play a substantial role in the evolution of the universe and large-scale structure formation. Measurements of baryon acoustic oscillations from the Dark Energy Spectroscopic Instrument (DESI) [6], in aggregate with observational data from Planck [7] and the Atacama Cosmology Telescope (ACT) [8], place the most stringent upper limit on the sum of neutrino mass eigenstates from cosmology: 72 meV with 95% confidence level [9]. Inherent to this bound is the assumption of the standard CDM cosmological framework [10], and the inclusion of model extensions or other nonstandard physics lessens the constraint ( [11, 12], for instance).
Additionally, neutrinos possess no electric or color charge, and have thus been posited to be their own antiparticle (a so-called “Majorana particle”) [13]. Experiments have sought to test this hypothesis by searching for neutrinoless double- decays [14] from various radioactive isotopes (such as [15, 16], [17], [18, 19], [20], and [21, 22], noninclusive). Outcomes of these searches set an upper limit window on the effective Majorana mass .333 . The smallest of these windows was set by KamLAND-Zen [21] at - meV, where the range is resultant from different nuclear calculations.
A more direct (and model-independent) method to measure the neutrino mass is to study the kinematics of electron capture [23] or -decay as originally proposed by Enrico Fermi [24]. The endpoint of the electron differential energy spectrum (i.e., maximal energy) in both cases is distorted directly according to the effective neutrino mass . While there are multiple ongoing efforts to study the electron capture kinematics of [25, 26], -decay experiments have thus far proven to be the more prolific approach.
Over the last 30 years, molecular tritium -decay experiments have been solely responsible for shrinking the world-leading upper limit on [27]. The current limit of < 450 meV with 90% confidence level was recently set by the Karlsruhe Tritium Neutrino (KATRIN) experiment [28]. The design of KATRIN, however, limits it to an idealized sensitivity of 200 meV [29], with background from surface radioimpurities [30], statistical uncertainty, and other systematic effects possibly resulting in a less-stringent final result [28]. As a whole, molecular tritium experiments are inherently limited to a sensitivity of 100 meV due to broadening caused by internal molecular motion [31, 32]. Instead using atomic tritium , recently proposed projects such as PTOLEMY suggest that advances in our understanding of neutrino physics and techniques for high-sensitivity instrumentation could significantly improve sensitivity to measuring [33, 34].
@@ 1.2 The PTOLEMY Experiment
PTOLEMY aims to create a scalable neutrino mass measurement experiment that performs precision spectroscopy of electrons produced near the endpoint energy from the -decay of atomic tritium: . With strong background suppression and high energy resolution, distortions of the emission spectrum endpoint would enable a direct measurement of . To achieve this, PTOLEMY proposes the following sequence: (1) guide tritium-emitted electrons into a uniform 1 T magnetic field region, (2) transmit them slowly through this uniform region via drift (in the meanwhile using the emitted cyclotron radiation to estimate particle kinematics and select near-endpoint candidates), (3) transport selected candidates out of the uniform region and drastically reduce their kinetic energy by counterbalancing motion from and (where is set according to the previously estimated kinematics), (4) and finally perform a highly precise measurement of the remaining kinetic energy.444For a particle exposed to an electric field while undergoing cyclotron motion due to a magnetic field (satisfying ), the drift velocity of its guiding center (the center point of circular cyclotron motion) is given as and commonly referred to as drift. For a non-uniform magnetic field, the total net force is given as , and a similar drift term arises: , where is the orbital magnetic moment, is charge, and is the kinetic energy component of the particle transverse to the magnetic field direction. A schematic of the PTOLEMY prototype is shown in Figure 1.
@@
Tritium was selected for multiple reasons, including suitable half-life ( years) and low -value ( keV) [35]. The proposed target substrate is graphene [36], which can efficiently store atomic tritium by binding it to carbon atoms. Extensive work is currently being undertaken by the collaboration and community at large to study the solid state effects on the emission spectra [37, 38, 39, 40, 41].
@@
To attain high measurement resolution, PTOLEMY utilizes transition-edge sensor (TES) microcalorimeters, which precisely measure energy via the strongly temperature-dependent resistance of a cryogenic superconducting material’s phase transition. In order for tritium endpoint electrons to fall within the dynamical range of the TES, static electromagnetic fields are used to filter their kinetic energy from 18.6 keV to roughly 10 eV. This is performed by PTOLEMY’s novel transverse drift filter [42]. The filter uniquely orients both the magnetic field gradient and electric field orthogonal to the magnetic field, resulting in components of the electron’s motion driven by and . The force produced by , when accompanied by E B drift, can do work on an electron and reduce its kinetic energy transverse to the magnetic field (). Namely, , indicating a change in voltage potential corresponds to a change in kinetic energy. Hence, a particle steadily loses kinetic energy as it climbs a higher potential barrier through the filter and travels towards the TES. Applying an additional electric field parallel to the magnetic field facilitates a reduction in parallel kinetic energy () and thus total kinetic energy (). The design of the transverse filter allows for a compact, scalable detector, with each dimension on the order of 1 meter [43]. In total, this technique could achieve an energy resolution as small as 50 meV [44], opening the possibility to a detection of the Cosmic Neutrino Background [45, 35] in addition to an effective neutrino mass measurement.
The drawback, however, is that a particle’s emission angle must be accurately known to properly tailor an orthogonal electric field in the filter that depletes (i.e., while near-endpoint electrons have similar energies, their momenta split with respect to the magnetic field can vary drastically). Further, this field necessarily must be set prior to the particle entering the filter to maintain electrostatic conditions. To accurately infer such emission spectra, PTOLEMY utilizes cyclotron radiation emission spectroscopy (CRES), which is a technique pioneered by the Project 8 collaboration [46, 47]. Using radio-frequency (RF) antennae, an electron’s total and perpendicular kinetic energy are estimated by the relativistic frequency shift of the emitted cyclotron radiation as it transits through a uniform 1 T magnetic field “RF region” (preceding the transverse drift filter) via drift. The RF region consists of electrostatic mirrors parallel to the magnet pole faces (to trap the particle) and an orthogonal electric field set according to voltage potentials. Particularly, drift is precisely slowed to trap electrons in the region long enough to obtain an adequate integrated radiated power signal for the CRES measurement. In the instance of an inferred near-endpoint electron, the estimate is used to adjust voltage potentials, producing a suitable electric field in the drift filter. Provided enough time to obtain a large cyclotron radiation signal, CRES has been shown to stringently limit backgrounds near the tritium endpoint [48], enabling the possibility of direct differential energy measurement in time coincidence following a non-destructive RF tag. With an energy resolution of , Project 8 has utilized CRES to establish a neutrino mass sensitivity of 155 eV [48]. Addressing plans and challenges to scale up the experiment and implement an atomic tritium source, they propose to achieve a sensitivity of 40 meV [49]. Instead, PTOLEMY uses CRES solely to obtain an estimate of the particle’s kinematics with substantially coarser resolution than the final TES measurement.
The collaboration has performed CRES measurements of conversion electrons emitted from gas contained in an electromagnetic trap, successfully identifying signal tracks akin to observations by Project 8 [50]; a publication of these results is expected soon. In addition, detailed simulation work has been undertaken to understand RF emission and antenna response in a system similar to the future RF region, including estimates of required electron time of flight and the development of a fitting algorithm to extract kinematic parameters [51]. However, electron transport as outlined for the RF region has yet to be shown in a physical setup. This is a crucial milestone in actualizing the detector, enabling future CRES measurements, filter implementation, and source injection.
@@ 1.3 Transport
drift has been extensively researched in the contexts of plasma dynamics, propulsion, and astrophysical phenomena (for instance: [52], [53], and [54]), though is seldom employed deliberately for charged particle transport. Notwithstanding, it is relatively straightforward to produce and implement, especially in setups already involving magnetic fields. The addition of electrodes in a Penning ion source test stand has been presented as a means to trap secondary electrons by E B drift [55], and KATRIN has shown the use of non-parallel electric and magnetic fields at the emission point of a photoelectron gun yields an apt angular-selective calibration source for their experiment [56]. Further yet, E B drift, in conjunction with an electrostatic barrier (i.e., the concepts utilized for transport in PTOLEMY’s RF region), has been implemented for positron injection into a magnetic dipole trap [57]. These prior studies, however, have no consideration of particle time of flight with respect to E B drift, nor do they manipulate electric fields to slow transit speed.555By time of flight, we are referring to the amount of time a particle takes to travel through a defined distance or volume. In this regard, PTOLEMY’s concept of slow drift for low-energy electron transport is entirely novel. Tethered to the strength of , an electron’s drift velocity can easily be changed by adjusting voltage potentials throughout the system. This feature is crucial to precisely reduce drift velocity near the RF antennas such that there is enough time (on the order of 100 s [51]) to obtain a large enough signal to accurately tag electrons.
@@ 2 Demonstrator Design
2.1 Aims and Scope
This paper presents a working physical setup that transports -decay electrons as described for the PTOLEMY RF region. Namely, an emitted electron is maintained within a defined volume via electrostatic mirrors and transits by E B drift. In the central portion of this volume, drift is reduced to mimic an injected particle entering the RF region, slowly passing through, then subsequently regaining drift speed as it approaches the filter. CRES measurements are not performed here; rather, transported electrons are determined according to a detector placed on the other side of the slow drift region. The recorded information is used to tabulate transmission rates.
Though we are primarily interested in maximally slowing the electron in the physical setup, a more detailed characterization is necessary to undergird a correspondence to simulation, thereby validating its function. With a programmable high voltage power supply, we record event rates under varied voltage potential configurations. We thus develop frameworks to systematically alter voltages and create coherent event rate datasets, which we call ‘Modes of Operation’ and describe further in Section 3.3. A correspondence between these datasets and their simulation counterparts confirms system function and allows us to extract time of flight information.
Importantly, this work intends to show the feasibility of actualizing slowed E B drift, including practices to facilitate the RF tagging capability. It does not serve as a high-precision instrument or is reflective of the future capabilities of the PTOLEMY RF region, but rather exhibits a novel technique in a simple setup. We conclude this section by discussing some perceived limitations.
1.
Attainable Time of Flight: To emulate the PTOLEMY RF region, we wish to observe slowed drift in a uniform magnetic field. Unfortunately, the uniform field volume of the electromagnet available for this study is fairly limited . This directly restricts the dimensions of the setup and, hence, electron time of flight. The situation is exacerbated due to a persistent residual electric field in the slow drift region (produced from other voltage potentials), which results in a sizable finite minimum field. We are additionally constrained (for reasons discussed in Section 2.7) to operate at a magnetic field of only 160 mT. In conjunction, this places a strict lower bound on achievable drift speed.
Though we show an manipulation of drift speed, the slowest attainable time of flight remains orders of magnitude smaller than the future needs of the detector ( ns, as compared to the previously quoted 100 s). We are confident, however, that significantly slower drift can be achieved with a larger setup and stronger magnetic field, enabling an electron containment of 100 s if not longer.
2.
Electron Source: For evident reasons of cost and handling, we choose to use a different radioactive isotope as a electron source in place of . We selected a source for its comparable -value, vacuum hardiness, and availability at our institution.
3.
Electron Detection: The primary function of a detector for this setup is to properly detect the transmitted spectrum above a minimum-energy threshold. A precise, high-resolution instrument tailored for low-energy electron detection would be superfluous for our needs. Under additional cost, size, and field resilience considerations, we opt for a commercial photodetector. With slight alterations, it is sensitive to a significant portion of the spectrum. Details are provided in Section 2.5.
The detector is intended to serve as a broad tag for electrons, in which case details of resolution and systematics are less important for the scope of this study. A comprehensive energy measurement would require full calibration, establishing an electron energy resolution (preferably with multiple mono-energetic sources), an evaluation of dead layer thickness (not provided by the manufacturer) and resultant average energy loss based on particle impingement angle, a meticulous assessment of the minor output changes under varied electric and magnetic fields, an understanding of changing output response due to self-heating, an estimation of electron energy loss due to cyclotron radiation emission and imperfect vacuum, and other possible systematic effects (such as noise contributions from other devices). While KATRIN has successfully characterized and employed a very similar detector for high-resolution measurements [58], the sole requirement for our system is to determine a minimum energy detector sensitivity (to properly compare with simulation). We resolve this by using a beta spectrometer, as discussed in the Supplemtary Material. With no available information regarding energy resolution, we approximate it with the manufacturer-provided photocurrent uncertainty (11%).
4.
Systematic Uncertainties: We are confident in system operation by identifying corresponding, broad trends between recorded data and simulation (manifesting as a least-squares regression fit). While important, a rigorous examination of systematic uncertainties is not necessary to achieve the outlined goal of this demonstration: feasibility, not precision. For consideration of future implementation, we discuss various contributions to uncertainty. However, we estimate uncertainty by assessing the fluctuation of repeated measurements, which we consider suitable for our purposes.
@@ 2.2 Concept
The conceptual design for the demonstrator is shown in Figure 2. Eight electrode plates are situated between the pole faces of a magnet. Two of these electrodes, the bouncing electrodes, are oriented parallel to the magnet pole faces and supplied with a negative voltage to create a potential barrier to contain electrons within the apparatus (such that the electron motion parallel to the magnetic field exhibits a bouncing behavior), similar to the electrostatic mirror described in [59]. The remaining six are organized into three pairs of vertically offset drift electrodes, with voltages applied to generate an electric field orthogonal to the magnetic field. The potential difference between the center pair of plates is set to be smaller than that of the outer pairs, thereby creating a region of slower drift. In total, an electron injected through one end of the demonstrator undergoes cyclotron motion, bounces between bouncing electrodes, and drifts toward the other end. An example of such a trajectory is shown from a top-down orientation in Figure 3. The increased density of trajectory lines in the center region indicates slower drift. No work is performed by drift, hence electron kinetic energy only deviates in accordance with the potential experienced from the bouncing electrodes throughout the trajectory. As with the physical setup, the detection location in this simulation is maintained at a voltage potential comparable to the emission point ( V), ensuring that the final electron kinetic energy roughly coincides with that originally emitted. This Figure and all subsequent simulation estimates are generated using CST Studio Suite by Dassault Systèmes [60].
@@ 2.5 PIN Diode Detector
For electron identification, the system size and operating environment necessitate the use of a detector that is both compact and resilient to high magnetic fields. Inspired by the utilization of PIN diodes for low-energy electron detection by KATRIN [58], the OSRAM BPX 61 Silicon PIN diode was selected as a detector for this study. The specific model was chosen for its commercial availability and comparatively low cost. PIN diodes are electrical components containing an undoped intrinsic semiconductor region between a p-type and n-type semiconductor. When operated in reverse bias, they are sensitive to energy depositions of charged particles and light. The BPX 61 additionally features a low dark current of 2 nA and a relatively large active detection area (), limiting electrical noise while increasing possible event count. However, commercial diodes — including the BPX 61 — are typically hermetically sealed by plastic or glass to prevent damage to the semiconductor region. Such covers, though providing little impedance to light, have a great enough stopping power to entirely prevent low-energy electron detection. For our study, the cover was removed and the exterior casing was flattened to the base of the diode, improving sensitivity to higher-energy electrons with larger cyclotron radii. A heat shrink tube was attached around the neck of the diode to prevent torsion/bending from magnetization. An image of the BPX 61 diode and an example of the performed alterations are shown in Figures 7.
Energy threshold characterization of the device was performed using a beta spectrometer (details provided in Supplementary Material). In summary, the spectrometer uses a solenoidal magnetic lensing system to generate a cylindrically-symmetric magnetic field. Emitted electrons undergoing a helical trajectory require a certain momentum (and hence energy) to pass circular barriers inside the device before reaching the detector. The spectrometer sequentially selects ranges of allowable momenta with a constant resolution of 2.45% () by incrementally increasing the strength of the solenoidal magnetic field. Converting momentum to energy, each increment corresponds to a specific source energy range where events are recorded by the PIN diode. Each setting lasts roughly 1 minute, and the full range of settings is cycled through on the order of 50 times (to account for time-dependent noise). Events are considered to be waveforms whose peak height exceeds a value of 45 mV above baseline after amplification (the electronic and amplification chain are identical to those discussed in Section 2.6). Based on the phase space of allowed particle trajectories in the system (discussed in Section 5), we situate the diode at a shallow angle () with respect to the instrument aperture. In Figure 8, we present a trigger efficiency curve for the device obtained from beta spectrometer counts, observing a half-maximal trigger count occurring at keV. This curve is then scaled to obtain a relative trigger efficiency. To clarify, this result does not reflect the true sensitivity of the device, but rather the trigger efficiency at a threshold safely above the noise floor of our subsequent data-taking conditions. Following manufacturer specifications on photocurrent uncertainty, we assign an energy resolution of 11%.